OpenStudy (anonymous):

What percent of 110 is 99?

OpenStudy (anonymous):

99 / 110 * 100 = 90 per cent

OpenStudy (anonymous):

99/110 = 9/ 10 (dividing top,bottom by 10) x by 100 = 900/10 = 90 %

OpenStudy (anonymous):

thank you